# Jian-hao Kanga, Zhun Goub and Nan-jing Huangc

### v1

@@ 3 Problem formulation and verification theorem
$\widehat{V}(t,x,m_{1},m_{2}):=\sup\limits_{u}\;\mathbb{E}_{t}\left[\varphi\left(X^{u}(T)+\beta M_{1}^{u}(T)\right)\right],$ (3.4)
@@
$\widehat{J}^{u}(t,x,m_{1},m_{2}):=\varphi^{-1}\left(\mathbb{E}_{t}\left[\varphi\left(X^{u}(T)+\beta M_{1}^{u}(T)\right)\right]\right),$ (3.5)
@@
$J^{u}(t,x,m_{1},m_{2}):=\int(\varphi^{\gamma})^{-1}\left(\mathbb{E}_{t}\left[\varphi^{\gamma}\left(X^{u}(T)+\beta M_{1}^{u}(T)\right)\right]\right)\mathrm{d}\Gamma(\gamma).$ (3.6)
@@
$y^{u,\gamma}(t,x,m_{1},m_{2}):=\mathbb{E}_{t}\left[\varphi^{\gamma}\left(X^{u}(T)+\beta M_{1}^{u}(T)\right)\right]$ (3.7)
@@
$J^{u}(t,x,m_{1},m_{2})=\int(\varphi^{\gamma})^{-1}\left(y^{u,\gamma}(t,x,m_{1},m_{2})\right)\mathrm{d}\Gamma(\gamma)$ (3.8)
@@
From equations (3.7) and (3.8), we observe that the equation (3.8) can be regarded as a summation of nonlinear functions dependent on expectations, leading to time-inconsistency in the optimization problem. To address the time-inconsistency issue, we build upon the studies of [9, 10, 17, 25, 39] to pursue the time-consistent equilibrium strategy for reinsurance and investment by employing a game-theoretic framework. To begin with, letting $\mathcal{Q}:=[0,T]\times\mathcal{O}$ with $\mathcal{O}:=\mathbb{R}\times\mathbb{R}\times\mathbb{R}$, we formally define the class of admissible reinsurance and investment strategies as follows.
@@
    for each initial point $(t,x,m_{1},m_{2})\in\mathcal{Q}$, the SDDE (2.3) possesses a unique strong solution $X^{u}$;
@@
    $\{u(t)\}_{t\in[0,T]}$ is $\mathcal{F}_{t}$-progressively measurable and continuous, $\mathbb{E}\left[\int^{T}_{0}\left(q^{2}(t)+\pi^{2}(t)(X^{u}(t))^{2}\right)\mathrm{d}t\right]<\infty$ and $q(t)\geq 0$;
@@
Then following game-theoretic approaches, we proceed to define the equilibrium reinsurance and investment strategy.
@@
For an arbitrary initial point $(t,x,m_{1},m_{2})\in\mathcal{Q}$, denote an admissible reinsurance and investment strategy by $\hat{u}(t,x,m_{1},m_{2})=(\hat{q}(t,x,m_{1},m_{2}),\hat{\pi}(t,x,m_{1},m_{2}))$. Consider arbitrary parameters $q\geq 0$, $\pi\in\mathbb{R}$ and $\varepsilon>0$, then define a reinsurance and investment strategy through
$\left(q_{\varepsilon}(s,\widetilde{x},\widetilde{m}_{1},\widetilde{m}_{2}),\pi_{\varepsilon}(s,\widetilde{x},\widetilde{m}_{1},\widetilde{m}_{2})\right)=\begin{cases}(q,\pi),\quad\quad\quad\quad\quad\;\quad\quad\;\quad\quad\quad\quad\quad\text{for}\quad\,t\leq s<t+\varepsilon,\\ (\hat{q}(s,\widetilde{x},\widetilde{m}_{1},\widetilde{m}_{2}),\hat{\pi}(s,\widetilde{x},\widetilde{m}_{1},\widetilde{m}_{2})),\quad\text{for}\;\quad t+\varepsilon\leq s<T,\end{cases}$
where $(\widetilde{x},\widetilde{m}_{1},\widetilde{m}_{2})\in\mathcal{O}$. Let $u_{\varepsilon}(s,\widetilde{x},\widetilde{m}_{1},\widetilde{m}_{2}):=(q_{\varepsilon}(s,\widetilde{x},\widetilde{m}_{1},\widetilde{m}_{2}),\pi_{\varepsilon}(s,\widetilde{x},\widetilde{m}_{1},\widetilde{m}_{2}))$. If
$\underset{\varepsilon\rightarrow 0}{\liminf}\frac{J^{\hat{u}}(t,x,m_{1},m_{2})-J^{u_{\varepsilon}}(t,x,m_{1},m_{2})}{\varepsilon}\geq 0$
for all $(q,\pi)\in\Pi$, then $\hat{u}(t,x,m_{1},m_{2})$ is defined as an equilibrium reinsurance and investment strategy, with its associated equilibrium value function $V(t,x,m_{1},m_{2}):=J^{\hat{u}}(t,x,m_{1},m_{2})$.
Pursuant to Definition 3.2, the insurer seeks to maximize the functional (3.8) by determining the equilibrium reinsurance and investment strategy. To streamline our presentation, we establish the following notations. For $\mathcal{D}=[0,T]\times\mathbb{R}\times\mathbb{R}$, denote
$C^{1,2,1}(\mathcal{D})=$ $\left\{\phi(t,x,m_{1})|\phi(t,\cdot,\cdot)\;\text{is once continuously differentiable on}\;[0,T],\right$.
$\;\left.\phi(\cdot,x,\cdot)\;\text{is twice continuously differentiable on}\;\mathbb{R}\;\text{and}\right$.
$\;\left.\phi(\cdot,\cdot,m_{1})\;\text{is once continuously differentiable on}\;\mathbb{R}\right\}$
and for any $\phi(t,x,m_{1})\in C^{1,2,1}(\mathcal{D})$ and $u=(q,\pi)$, let
$\mathcal{A}^{u}\phi(t,x,m_{1})=$ $\phi_{t}(t,x,m_{1})+\left[(A+\pi(\mu-r))x+Bm_{1}+Cm_{2}+a\eta+a\eta_{2}q\right]\phi_{x}(t,x,m_{1})$
$+0.5\left(b^{2}q^{2}+\pi^{2}x^{2}\sigma^{2}\right)\phi_{xx}(t,x,m_{1})+\left(x-\alpha m_{1}-e^{-\alpha h}m_{2}\right)\phi_{m_{1}}(t,x,m_{1}).$
@@
    $\mathcal{A}^{u}\phi(t,X^{u}(t),M_{1}^{u}(t))\in\mathbb{L}^{1}_{\mathcal{F}}(0,T;\mathbb{R})$ and $bq(t)\phi_{x}(t,X^{u}(t),M_{1}^{u}(t)),\;\pi(t)X^{u}(t)\sigma\phi_{x}(t,X^{u}(t),M_{1}^{u}(t))\in\mathbb{L}^{2}_{\mathcal{F}}(0,T;\mathbb{R})$.
@@
If $G(t,x,m_{1})\in C^{1,2,1}(\mathcal{D})$ and $G(t,X^{u}(t),M_{1}^{u}(t))$ satisfies the condition $(A1)$, then
$\mathrm{d}G(t,x,m_{1})=\mathcal{A}^{u}G(t,x,m_{1})\mathrm{d}t+bqG_{x}(t,x,m_{1})\mathrm{d}W_{1}(t)+\pi x\sigma G_{x}(t,x,m_{1})\mathrm{d}W_{2}(t).$
@@
Chang et al. [15] and Shen and Zeng [30] noted that optimal control problems involving delays typically constitute infinite-dimensional problems. To obtain analytical optimal solutions, additional constraints need to be imposed. For instance, the value function $V$ is required to depend solely on $(t,x,m_{1})$. In fact, since $M_{1}^{u}(t)$ is a process correlated with $M_{2}^{u}(t)$, we will subsequently have to specify the exact conditions under which $V$ maintains dependence only on $(t,x,m_{1})$. We now establish a verification theorem to characterize the equilibrium strategy for reinsurance and investment.
@@
Suppose there exist functions $U,Y^{\gamma},H\in C^{1,2,1}(\mathcal{D})$ satisfying
$\sup\limits_{u}\;\left\{\mathcal{A}^{u}U(t,x,m_{1})-\mathcal{A}^{u}H(t,x,m_{1})+\int\iota^{\gamma}(Y^{\gamma}(t,x,m_{1}))\mathcal{A}^{u}Y^{\gamma}(t,x,m_{1})\mathrm{d}\Gamma(\gamma)\right\}=0$ (3.9)
@@
$\mathcal{A}^{\hat{u}}Y^{\gamma}(t,x,m_{1})=0$ (3.10)
with boundary conditions $U(T,x,m_{1})=x+\beta m_{1}$ and $Y^{\gamma}(T,x,m_{1})=\varphi^{\gamma}(x+\beta m_{1})$ for all $\gamma$, where
$H(t,x,m_{1})=\int(\varphi^{\gamma})^{-1}\left(Y^{\gamma}(t,x,m_{1})\right)\mathrm{d}\Gamma(\gamma)$ (3.11)
@@
$\hat{u}=\arg\sup\limits_{u}\;\left\{\mathcal{A}^{u}U(t,x,m_{1})-\mathcal{A}^{u}H(t,x,m_{1})+\int\iota^{\gamma}(Y^{\gamma}(t,x,m_{1}))\mathcal{A}^{u}Y^{\gamma}(t,x,m_{1})\mathrm{d}\Gamma(\gamma)\right\}.$
@@
In the proof of Theorem 3.1, we can obtain that $Y^{\gamma}(t,x,m_{1})=y^{\hat{u},\gamma}(t,x,m_{1},m_{2})$ and $U(t,x,m_{1})=J^{\hat{u}}(t,x,m_{1},m_{2})$ from equations (C.46) and (C). This implies that $y^{\hat{u},\gamma}$ and $J^{\hat{u}}$ are independent of $m_{2}$. In the sequel, we will adopt the notations $y^{\hat{u},\gamma}(t,x,m_{1})$ and $J^{\hat{u}}(t,x,m_{1})$ for simplicity.
Prior to concluding this section, we formulate a more explicit expression of the pseudo HJB equation to characterize the equilibrium strategy for reinsurance and investment. For this purpose, by using the equation (3.11), one can derive that
$\left\{\begin{aligned} &H_{t}(t,x,m_{1})=\int\iota^{\gamma}\left(Y^{\gamma}(t,x,m_{1})\right)Y_{t}^{\gamma}(t,x,m_{1})\mathrm{d}\Gamma(\gamma),\\ &H_{x}(t,x,m_{1})=\int\iota^{\gamma}(Y^{\gamma}(t,x,m_{1}))Y_{x}^{\gamma}(t,x,m_{1})\mathrm{d}\Gamma(\gamma),\\ &H_{m_{1}}(t,x,m_{1})=\int\iota^{\gamma}(Y^{\gamma}(t,x,m_{1}))Y_{m_{1}}^{\gamma}(t,x,m_{1})\mathrm{d}\Gamma(\gamma),\\ &H_{xx}(t,x,m_{1})=\int\iota^{\gamma}(Y^{\gamma}(t,x,m_{1}))Y_{xx}^{\gamma}(t,x,m_{1})\mathrm{d}\Gamma(\gamma)+\int(\iota^{\gamma})^{{}^{\prime}}(Y^{\gamma}(t,x,m_{1}))(Y_{x}^{\gamma}(t,x,m_{1}))^{2}\mathrm{d}\Gamma(\gamma),\end{aligned}\right$. (3.12)

### v2

@@ 3 Problem formulation and verification theorem
$\widehat{V}(t,x,s_{1},m_{1},m_{2}):=\sup\limits_{u}\;\mathbb{E}_{t}\left[\varphi\left(X^{u}(T)+\beta M_{1}^{u}(T)\right)\right],$ (3.4)
@@
$\widehat{J}^{u}(t,x,s_{1},m_{1},m_{2}):=\varphi^{-1}\left(\mathbb{E}_{t}\left[\varphi\left(X^{u}(T)+\beta M_{1}^{u}(T)\right)\right]\right),$ (3.5)
@@
$J^{u}(t,x,s_{1},m_{1},m_{2}):=\int(\varphi^{\gamma})^{-1}\left(\mathbb{E}_{t}\left[\varphi^{\gamma}\left(X^{u}(T)+\beta M_{1}^{u}(T)\right)\right]\right)\mathrm{d}\Gamma(\gamma).$ (3.6)
@@
$y^{u,\gamma}(t,x,s_{1},m_{1},m_{2}):=\mathbb{E}_{t}\left[\varphi^{\gamma}\left(X^{u}(T)+\beta M_{1}^{u}(T)\right)\right]$ (3.7)
@@
$J^{u}(t,x,s_{1},m_{1},m_{2})=\int(\varphi^{\gamma})^{-1}\left(y^{u,\gamma}(t,x,s_{1},m_{1},m_{2})\right)\mathrm{d}\Gamma(\gamma)$ (3.8)
@@
From equations (3.7) and (3.8), we observe that the equation (3.8) can be regarded as a summation of nonlinear functions dependent on expectations, leading to time-inconsistency in the optimization problem. To address the time-inconsistency issue, we build upon the studies of [9, 10, 17, 25, 38] to pursue the time-consistent equilibrium strategy for reinsurance and investment by employing a game-theoretic framework. To begin with, letting $\mathcal{Q}:=[0,T]\times\mathcal{O}$ with $\mathcal{O}:=\mathbb{R}\times[0,\infty)\times\mathbb{R}\times\mathbb{R}$, we formally define the class of admissible reinsurance and investment strategies as follows.
@@
    for each initial point $(t,x,s_{1},m_{1},m_{2})\in\mathcal{Q}$, the SDDE (2.3) possesses a unique strong solution $X^{u}$;
@@
    $\{u(t)\}_{t\in[0,T]}$ is $\mathcal{F}_{t}$-progressively measurable and continuous, $\mathbb{E}\left[\int^{T}_{0}\left(q^{2}(t)+\pi^{2}(t)(X^{u}(t))^{2}(S^{\delta}_{1}(t))^{2}\right)\mathrm{d}t\right]<\infty$ and $q(t)\geq 0$;
@@
Then following game-theoretic ways, we will define the equilibrium reinsurance and investment strategy.
@@
For an arbitrary initial point $(t,x,s_{1},m_{1},m_{2})\in\mathcal{Q}$, denote an admissible reinsurance and investment strategy by $\hat{u}(t,x,s_{1},m_{1},m_{2})=(\hat{q}(t,x,s_{1},m_{1},m_{2}),\hat{\pi}(t,x,s_{1},m_{1},m_{2}))$. Consider arbitrary parameters $q\geq 0$, $\pi\in\mathbb{R}$ and $\varepsilon>0$, then define a reinsurance and investment strategy through
$\left(q_{\varepsilon}(s,\widetilde{x},\tilde{s}_{1},\widetilde{m}_{1},\widetilde{m}_{2}),\pi_{\varepsilon}(s,\widetilde{x},\tilde{s}_{1},\widetilde{m}_{1},\widetilde{m}_{2})\right)=\begin{cases}(q,\pi),\quad\quad\quad\quad\quad\;\quad\quad\;\quad\quad\quad\quad\quad\quad\quad\;\;\text{for}\quad\,t\leq s<t+\varepsilon,\\ (\hat{q}(s,\widetilde{x},\tilde{s}_{1},\widetilde{m}_{1},\widetilde{m}_{2}),\hat{\pi}(s,\widetilde{x},\tilde{s}_{1},\widetilde{m}_{1},\widetilde{m}_{2})),\quad\text{for}\;\quad t+\varepsilon\leq s<T,\end{cases}$
where $(\widetilde{x},\tilde{s}_{1},\widetilde{m}_{1},\widetilde{m}_{2})\in\mathcal{O}$. Let $u_{\varepsilon}(s,\widetilde{x},\tilde{s}_{1},\widetilde{m}_{1},\widetilde{m}_{2}):=(q_{\varepsilon}(s,\widetilde{x},\tilde{s}_{1},\widetilde{m}_{1},\widetilde{m}_{2}),\pi_{\varepsilon}(s,\widetilde{x},\tilde{s}_{1},\widetilde{m}_{1},\widetilde{m}_{2}))$. If
$\underset{\varepsilon\rightarrow 0}{\liminf}\frac{J^{\hat{u}}(t,x,s_{1},m_{1},m_{2})-J^{u_{\varepsilon}}(t,x,s_{1},m_{1},m_{2})}{\varepsilon}\geq 0$
for all $(q,\pi)\in\Pi$, then $\hat{u}(t,x,s_{1},m_{1},m_{2})$ is defined as an equilibrium reinsurance and investment strategy, with its associated equilibrium value function $V(t,x,s_{1},m_{1},m_{2}):=J^{\hat{u}}(t,x,s_{1},m_{1},m_{2})$.
Pursuant to Definition 3.2, the insurer seeks to maximize the functional (3.8) by determining the equilibrium reinsurance and investment strategy. To streamline our presentation, we establish the following notations. For $\mathcal{D}=[0,T]\times\mathbb{R}\times[0,\infty)\times\mathbb{R}$, denote
$C^{1,2,2,1}(\mathcal{D})=$ $\left\{\phi(t,x,s_{1},m_{1})|\phi(t,\cdot,\cdot,\cdot)\;\text{is once continuously differentiable on}\;[0,T],\right$.
$\;\left.\phi(\cdot,x,\cdot,\cdot)\;\text{is twice continuously differentiable on}\;\mathbb{R},\;\right$.
$\;\left.\phi(\cdot,\cdot,s_{1},\cdot)\;\text{is twice continuously differentiable on}\;[0,\infty)\;\text{and}\right$.
$\;\left.\phi(\cdot,\cdot,\cdot,m_{1})\;\text{is once continuously differentiable on}\;\mathbb{R}\right\}$
and for any $\phi(t,x,s_{1},m_{1})\in C^{1,2,2,1}(\mathcal{D})$ and $u=(q,\pi)$, let
$\mathcal{A}^{u}\phi(t,x,s_{1},m_{1})=$ $\phi_{t}(t,x,s_{1},m_{1})+\left[(A+\pi(\mu-r))x+Bm_{1}+Cm_{2}+a\eta+a\eta_{2}q\right]\phi_{x}(t,x,s_{1},m_{1})$
$+0.5\left(b^{2}q^{2}+\pi^{2}x^{2}\sigma^{2}s^{2\delta}_{1}\right)\phi_{xx}(t,x,s_{1},m_{1})+\mu s_{1}\phi_{s_{1}}(t,x,s_{1},m_{1})+0.5\sigma^{2}s^{2\delta+2}_{1}\phi_{s_{1}s_{1}}(t,x,s_{1},m_{1})$
$+\pi x\sigma^{2}s^{2\delta+1}_{1}\phi_{xs_{1}}(t,x,s_{1},m_{1})+\left(x-\alpha m_{1}-e^{-\alpha h}m_{2}\right)\phi_{m_{1}}(t,x,s_{1},m_{1}).$
@@
    $\mathcal{A}^{u}\phi(t,X^{u}(t),S_{1}(t),M_{1}^{u}(t))\in\mathbb{L}^{1}_{\mathcal{F}}(0,T;\mathbb{R})$ and $\sigma S^{\delta+1}_{1}(t)\phi_{s_{1}}(t,X^{u}(t),S_{1}(t),M_{1}^{u}(t))$, $bq(t)\phi_{x}(t,X^{u}(t),S_{1}(t),M_{1}^{u}(t))$, $\pi(t)X^{u}(t)\sigma S^{\delta}_{1}(t)\phi_{x}(t,X^{u}(t),S_{1}(t),M_{1}^{u}(t))\in\mathbb{L}^{2}_{\mathcal{F}}(0,T;\mathbb{R})$.
@@
If $G(t,x,s_{1},m_{1})\in C^{1,2,2,1}(\mathcal{D})$ and $G(t,X^{u}(t),S_{1}(t),M_{1}^{u}(t))$ satisfies the condition $(A1)$, then
$\mathrm{d}G(t,x,s_{1},m_{1})=$ $\mathcal{A}^{u}G(t,x,s_{1},m_{1})\mathrm{d}t+bqG_{x}(t,x,s_{1},m_{1})\mathrm{d}W_{1}(t)$
$+\left[\sigma s^{\delta+1}_{1}G_{s_{1}}(t,x,s_{1},m_{1})+\pi x\sigma s^{\delta}_{1}G_{x}(t,x,s_{1},m_{1})\right]\mathrm{d}W_{2}(t).$
@@
We now establish a verification theorem to characterize the equilibrium strategy for reinsurance and investment.
@@
Suppose there exist functions $U,Y^{\gamma},H\in C^{1,2,2,1}(\mathcal{D})$ satisfying
$\sup\limits_{u}\;\left\{\mathcal{A}^{u}U(t,x,s_{1},m_{1})-\mathcal{A}^{u}H(t,x,s_{1},m_{1})+\int\iota^{\gamma}(Y^{\gamma}(t,x,s_{1},m_{1}))\mathcal{A}^{u}Y^{\gamma}(t,x,s_{1},m_{1})\mathrm{d}\Gamma(\gamma)\right\}=0$ (3.9)
@@
$\mathcal{A}^{\hat{u}}Y^{\gamma}(t,x,s_{1},m_{1})=0$ (3.10)
with boundary conditions $U(T,x,s_{1},m_{1})=x+\beta m_{1}$ and $Y^{\gamma}(T,x,s_{1},m_{1})=\varphi^{\gamma}(x+\beta m_{1})$ for all $\gamma$, where
$H(t,x,s_{1},m_{1})=\int(\varphi^{\gamma})^{-1}\left(Y^{\gamma}(t,x,s_{1},m_{1})\right)\mathrm{d}\Gamma(\gamma)$ (3.11)
@@
$\hat{u}=\arg\sup\limits_{u}\;\left\{\mathcal{A}^{u}U(t,x,s_{1},m_{1})-\mathcal{A}^{u}H(t,x,s_{1},m_{1})+\int\iota^{\gamma}(Y^{\gamma}(t,x,s_{1},m_{1}))\mathcal{A}^{u}Y^{\gamma}(t,x,s_{1},m_{1})\mathrm{d}\Gamma(\gamma)\right\}.$
@@
It follows from the proof of Theorem 3.1 that $Y^{\gamma}(t,x,s_{1},m_{1})=y^{\hat{u},\gamma}(t,x,s_{1},m_{1},m_{2})$ and $U(t,x,s_{1},m_{1})=J^{\hat{u}}(t,x,s_{1},m_{1},m_{2})=V(t,x,s_{1},m_{1},m_{2})$. This implies $y^{\hat{u},\gamma}$, $J^{\hat{u}}$ and $V$ are independent of $m_{2}$. In the sequel, we will adopt notations $y^{\hat{u},\gamma}(t,x,s_{1},m_{1})$, $J^{\hat{u}}(t,x,s_{1},m_{1})$ and $V(t,x,s_{1},m_{1})$ for simplicity.
4 Equilibrium strategies under the CEV model
In this section, we first seek to tackle the optimization problem of delayed reinsurance and investment with random risk aversion under the CEV model, adopting a general utility function framework to the greatest extent possible. Afterward, to derive (semi-)analytical equilibrium strategy for reinsurance and investment, we narrow our scope to analyze the exponential utility function.
Following [17, 25], one can formulate a more explicit expression of the pseudo HJB equation to characterize the equilibrium strategy for reinsurance and investment. For this purpose, by the equation (3.11), one has
$\left\{\begin{aligned} H_{t}(t,x,s_{1},m_{1})=&\int\iota^{\gamma}\left(Y^{\gamma}(t,x,s_{1},m_{1})\right)Y_{t}^{\gamma}(t,x,s_{1},m_{1})\mathrm{d}\Gamma(\gamma),\\ H_{x}(t,x,s_{1},m_{1})=&\int\iota^{\gamma}(Y^{\gamma}(t,x,s_{1},m_{1}))Y_{x}^{\gamma}(t,x,s_{1},m_{1})\mathrm{d}\Gamma(\gamma),\\ H_{s_{1}}(t,x,s_{1},m_{1})=&\int\iota^{\gamma}(Y^{\gamma}(t,x,s_{1},m_{1}))Y_{s_{1}}^{\gamma}(t,x,s_{1},m_{1})\mathrm{d}\Gamma(\gamma),\\ H_{m_{1}}(t,x,s_{1},m_{1})=&\int\iota^{\gamma}(Y^{\gamma}(t,x,s_{1},m_{1}))Y_{m_{1}}^{\gamma}(t,x,s_{1},m_{1})\mathrm{d}\Gamma(\gamma),\\ H_{xx}(t,x,s_{1},m_{1})=&\int\iota^{\gamma}(Y^{\gamma}(t,x,s_{1},m_{1}))Y_{xx}^{\gamma}(t,x,s_{1},m_{1})\mathrm{d}\Gamma(\gamma)\\ &+\int(\iota^{\gamma})^{{}^{\prime}}(Y^{\gamma}(t,x,s_{1},m_{1}))(Y_{x}^{\gamma}(t,x,s_{1},m_{1}))^{2}\mathrm{d}\Gamma(\gamma),\\ H_{s_{1}s_{1}}(t,x,s_{1},m_{1})=&\int\iota^{\gamma}(Y^{\gamma}(t,x,s_{1},m_{1}))Y_{s_{1}s_{1}}^{\gamma}(t,x,s_{1},m_{1})\mathrm{d}\Gamma(\gamma)\\ &+\int(\iota^{\gamma})^{{}^{\prime}}(Y^{\gamma}(t,x,s_{1},m_{1}))(Y_{s_{1}}^{\gamma}(t,x,s_{1},m_{1}))^{2}\mathrm{d}\Gamma(\gamma),\\ H_{xs_{1}}(t,x,s_{1},m_{1})=&\int\iota^{\gamma}(Y^{\gamma}(t,x,s_{1},m_{1}))Y_{xs_{1}}^{\gamma}(t,x,s_{1},m_{1})\mathrm{d}\Gamma(\gamma)\\ &+\int(\iota^{\gamma})^{{}^{\prime}}(Y^{\gamma}(t,x,s_{1},m_{1}))Y_{x}^{\gamma}(t,x,s_{1},m_{1})Y_{s_{1}}^{\gamma}(t,x,s_{1},m_{1})\mathrm{d}\Gamma(\gamma),\end{aligned}\right$. (4.12)